OpenStudy (godlovesme):

help plz i'll give medals for correct answers :) thanks! <3

OpenStudy (godlovesme):

Angela and Neil are going to the movies. They each bought a medium popcorn, and Neil got a small soft drink. Angela had a $5 gift certificate to put toward the cost, and Neil paid the rest, which came to $28.80. A movie ticket costs $10.50 and a medium popcorn costs $5.10. How much does a small soft drink cost at the theater?

OpenStudy (godlovesme):

a. $2.60 b. $2.80 c. $18.20 d. $7.70

OpenStudy (godlovesme):

@Nnesha @Loser66 @Michele_Laino please help

Directrix (directrix):

Popcorn: 2 * 5.10 Drink: x Tickets: 2 * 10.50 Coupon: $5 Paid by Neil: 28.80

Directrix (directrix):

2 * 5.10 + x + 2 * 10.50 - 5 = 28.80

Directrix (directrix):

Solve for x.

OpenStudy (godlovesme):

i'll hang on @Directrix

Directrix (directrix):

Okay @Godlovesme We have to solve this for x: 2 * 5.10 + x + 2 * 10.50 - 5 = 28.80

OpenStudy (anonymous):

Look at the chat bar for Math. I gave a long explanation, and for some reason I could not post it.

Directrix (directrix):

Take the price that Neil paid and add the 5 dollar certificate in order to get a total cost of all the concessions. You should come up with a grand total of 33.80. Obviously they both got movie tickets, so subtract 21 dollars from the 33.80 total. You should come up with 12.80. Next, take the cost of both popcors (which would be 10.20 and subtract that form the 12.80 left. you should come up with 2.60 as the remaining. So, the cost of the small drink is 2.60. This is the work of @csheveland

OpenStudy (godlovesme):

2*5.10 + x + 2*10.50 -5 = 28.80 10.2 + x + 21 - 5 = 28.80 x= 28.80 - 26.20 x=2.6 :D @directrix

OpenStudy (anonymous):

All you have to do when solving the equation that is solve for x. Thanks man I appreciate it.

Directrix (directrix):

Just as @cleveland said.

OpenStudy (anonymous):

You reply was excellente.

Directrix (directrix):

@csheveland You are welcome. This website does some crazy things.

Directrix (directrix):

Whatever this forum is --> UPM F��sica de Bachillerato I am stuck in there and cannot get out.

OpenStudy (anonymous):

Lol.

Directrix (directrix):

I went to the forum to see what was there. Nothing - no questions, no answers. So, I clicked to get out and am stuck there. At least I can get to math.

OpenStudy (anonymous):

Um... You can't even log out?

OpenStudy (godlovesme):

thanks so much guys @Directrix @csheveland :) idk who to medal tho lol

Directrix (directrix):

There are 3 of us and we can each give a medal. So, we are in good shape except for you, @Godlovesme . I have already given my medal.

OpenStudy (anonymous):

Did I give a medal yet? I don't really btw, although I love my biology medals.

Directrix (directrix):

>>Um... You can't even log out? @csheveland No. On the forum list, it does not show up as my studying there. On the math bar, I can see it listed but the x will not x out. So, I am stuck with that chat room. I would say go there and see what you think but then you'd be stuck with it, too. At least, I can write on the threads, so far. lol

OpenStudy (anonymous):

lol. ok i think im in...

OpenStudy (anonymous):

I saw you in Biology section.